13+5x-2x=37 what does x equal show work thanks
Answer by AnlytcPhil(1806)

13+5x-2x=37 what does x equal show work thanks


13 + 5x - 2x = 37

First combine the like terms on the left which are
+5x and -2x.  That gives you +3x. So you replace
"+ 5x - 2x" by "+ 3x". So now you have this:

    13 + 3x = 37 

Since you want to get the x-term, +3x, all by itself
we get rid of the 13 term on the left by adding its
opposite, -13, to both sides

    13 + 3x = 37
   -13       -13
------------------
         3x = 24

When we get to this stage, where there is just one 
term on the left consisting of a coefficient and the
unknown, then you divide both sides by that 
coefficient, which is 3.
 
         3x     24
        ---- = ----
          3      3

Then the 3's cancel on the left and the 3 goes into the
24 8 times on the right:

         1       8
         3x     24
        ---- = ----
          3      3
          1      1

so the answer is:

           x = 8

Edwin